What volume of 18.0 mol/L sulfuric acid must be used to prepare 15.5 L of 0.178 mol/L H2SO4
Answer #1

Here 18.0 mol/L sulfuric acid is being diluted to a 0.178 mol/L solution.

The volume (V1) of a solution of concentration M1 required to dilute the solution to a concentration of M2 and volume V2 is given by:

V_1=\frac{M_2V_2}{M_1}

Here we have:

M1 = 18.0 mol/L

M2 = 0.178 mol/L

V2 = 15.5 L

Therefore the volume of 18.0 mol/L sulfuric acid to be used is:

V_1=\frac{M_2V_2}{M_1}

V_1=\frac{0.178\: \: mol/L\times 15.5\: \: L}{18.0\: \: mol/L}=0.153\: \: L

Hence the volume of 18.0 mol/L sulfuric acid to be used is 0.153 L or 153 mL.